Verify that the indicated family of functions is a solution of the given differential equation. Assume an appropriate intervaPPlease answer using MATLAB
Answer #1
The given diffential equation is 
\frac{d^2y}{dx^2}-6\frac{dy}{dx}+9y=0,   y=c_1e^{3x}+c_2xe^{3x}.

\text{Now we find}\quad \frac{dy}{dx} \quad \text{and} \quad \frac{d^2y}{dx^2}\quad \text{of}\quad y=c_1e^{3x}+c_2xe^{3x},\quad \text{by using matlab}Matlab code single derivative
syms x c1 c2
diff(c1*exp(3x)+c2*x*exp(3x),x)

Ans=

3*c1*exp(3x)+c2*exp(3x)+3*c2*x*exp(3x)

that is

\frac{dy}{dx}=3c_1e^{3x}+c_2e^{3x}+3c_2xe^{3x}

Matlab code double derivative

syms x c1 c2
diff(diff(c1*exp(3x)+c2*x*exp(3x),x))

Ans=

9*c1*exp(3x)+6*c2*exp(3x)+9*c2*x*exp(3x)

that is

\frac{d^2y}{dx^2}=9c_1e^{3x}+6c_2e^{3x}+9c_2xe^{3x}

The given differential equation in term of x, is

\frac{d^2y}{dx^2}-6 \frac{dy}{dx}+9y=9c_1e^{3x}+6c_2e^{3x}+9c_2xe^{3x}-6(3c_1e^{3x}+c_2e^{3x}+3c_2xe^{3x})+{\color{Red} 9(c_1e^{3x}+c_2xe^{3x})}\\=0